What is the equation of the line that has a slope of -2 and passes (3, -4)?

A. y - 4= -2x + 3
B. y - 4= -2(x + 3)
C. y+4 = -2(x – 3)
D. y = -2x -12

Answer by nerdybill(7384)

What is the equation of the line that has a slope of -2 and passes (3, -4)?
Basically, you would use the "point-slope" form:
y - y1 = m(x-x1)
Doing so gives you:
y - (-4) = -2(x-3)
Which is:
.
C. y+4 = -2(x – 3)